# Robust and Flexible Omnidirectional Depth Estimation with Multiple 360 Cameras

### v1

@@ IV-B1 Pairwise Stereo Matching
IV-B2 Omnidirectional Depth Fusion
@@
We align the depth maps and confidence maps to the same viewpoint based on the extrinsic matrix and visibility. As shown in Figure 5, the depth fusion network generally follows Unet[64], containing two encoders and one decoder. One encoder takes concatenation of the aligned depth maps and confidence maps as input to effectively aggregate the depth feature. and the other takes RGB panoramas as input to extract context and boundary features. Subsequently, these two types of features are fused through a multi-scale feature fusion block to generate the more comprehensive and informative feature maps. Finally, the decoder utilizes fused feature maps to perform regression and predict the final depth map.
@@
IV-C Spherical Sweeping Multi-view Omnidirectional Depth Estimation(SSMODE)
@@
As shown in Figure 6, the proposed SSMODE first extracts features of input panoramas with the Spherical Feature Extraction Module. By employing the GEER projection, we establish the angular coordinate system $(\theta,\phi)$ and define a collection of hypothetical spheres at different depths $\rho_{i}$. The features for each point $(\rho_{i},\theta,\phi)$ across different views are obtained through the camera extrinsics, and these features are concatenated to construct the spherical cost volume. As illustrated in Figure 3 and Equation(5), the unified spherical cost volume contains the matching cost of each pixel at each hypothetical depth. Similar to stereo matching, the spherical cost volume is also represented as a 5D tensor with a shape of $(B\times C\times D\times H\times W)$, where $(H\times W)$ denotes the angular coordinate of the sphere and $D$ represetns the number of hypothetical spheres. Subsequently, the 3D stacked hourglass module is employed to aggregate the multi-view spherical matching cost. Based on the regressed comprehensive matching cost, the weights of different depth are calculated, and the final depth is obtained by weighted summation:
@@
V Dataset
TABLE I: Overview of the proposed datasets and other published datasets
@@ V Dataset
As summarized in II-C, although many datasets have been proposed for omnidirectional depth estimation, no 360 stereo dataset for outdoor road scenes is available due to the difficulty of acquiring 360 outdoor 3D datasets in the real world. Therefore, we create a public available 360 multi-view dataset Deep360 based on the CARLA autonomous driving simulator. Figure 7 shows some examples of the dataset. We set four 360°cameras and arrange the cameras on a horizontal plane to form a square with side length as one meter, as shown in Figure 1(a). The cameras are numbered from 1 to 4. Any two of the cameras can form a stereo pair, so there are 6 $(C_{4}^{2})$ pairs in total. Each frame consists of six pairs of rectified panoramas, which cover all the pairwise combinations of four 360 cameras, six corresponding disparity maps and one ground truth depth map. All these images and maps have a resolution of $1024\times 512$. To acquire realistic 360 outdoor road scenes with high variety, we make the car with 360 cameras in CARLA drive automatically [56] in six different towns and spawn many other random pedestrian and vehicles.
We also provide a soiled version of the Deep360 dataset, which can be used to train and evaluate 360 depth estimation algorithms under harsh circumstances in autonomous driving. Deep360-Soiled contains panoramas soiled or affected by three common outdoor factors: mud spots, water drops and glare, as illustrated in Figure 1(e). An overview of the proposed dataset and other published 360 datasets is listed in Table I.
VI Experiment Results
TABLE II: Quantitative results of stereo matching on the proposed Deep360 dataset. The metrics refer to disparity errors
@@ VI Experiment Results
VI-A Experiment Settings
VI-A1 Datasets
@@ VI-A3 Metrics
VI-B Omnidirectional Stereo Matching
The existing binocular stereo matching algorithm is able to directly predict the spherical binocular disparity map at arbitrary relative positions based on the GEER projection method. Thus, we first evaluate the proposed omnidirectional stereo matching network on the Deep360 dataset and compare it with the excellent stereo matching algorithms PSMNet[20], AANet [24] and CREStereo[31], as well as the omnidirectional method 360SD-Net[5]. For these approaches, we use the pre-trained models from the authors and follow their hyperparameters to finetune on Deep360. Figure 8 shows the qualitative results of omnidirectional stereo matching on the deep360 dataset. The quantitative results in Table II illustrate that our stereo matching network with spherical feature learning achieves SOTA performance on 360 stereo matching.
We also present the results of stereo matching of two 360 cameras at different relative positions in Figure 9. As shown in Figure 8 and Figure 9, the proposed GEER projection establishes the epipolar constraint of binocular 360 cameras at arbitrary relative positions. The results show that the proposed stereo matching method with spherical feature extraction module achieves high precision with clear details.
VI-C Omnidirectional Depth Estimation
@@
TABLE III: Quantitative results of omnidirectional depth estimation on the proposed Deep360 dataset. The metrics refer to depth errors
@@
As shown in Table III, PSMODE and SSMODE perform favorably against SOTA omnidirectional depth estimation methods, especially on the dataset with soiled panoramas. We also compare the result of PSMODE with and without the fusion stage in Table III. As the results show, the multi-view depth fusion stage significantly improves the accuracy of omnidirectional depth estimation. As demonstrated in Table III and Figure 11, the accuracy degradation of the proposed methods on the soiled data is significantly lower than that of existing methods. The comparison demonstrates the robustness of the proposed multi-view depth estimation methods against camera soiling. We also evaluate the proposed methods on 3D60 dataset and illustrate the results in Table IV and Figure 10. The proposed PSMODE and SSMODE achieve high accuracy on both indoor and outdoor scenes. In this paper, we leverage all three stereo pairs within the 3D60 (left-right,up-down,up-right) in the depth fusion stage of PSMODE by employing the GEER projection. Thus, the results in Table IV is better than those reported in the conference version[11].
TABLE IV: Quantitative results of omnidirectional depth estimation on 3D60 dataset. The metrics refer to depth errors
@@
VI-D Results on Real Scenes
We use the best PSMODE model trained on Deep360 to predict 360 depth maps on real-scene data. We use four Insta One X2 360 cameras to build the camera system, as shown in Figure 1(b). Figure 12 illustrates that the proposed algorithm also achieves an accurate depth estimation on real-scene data.
@@ VI-D Results on Real Scenes
VI-E Evaluation of Different Numbers of Views
@@
To evaluate the performance of PSMODE and SSMODE under varying view conditions, we conducted experiments on the clear and soiled Deep360 dataset using different numbers of views (4, 3, 2). Table V and Figure 13 indicate that as the number of views decreases, both PSMODE and SSMODE experience increase in error of depth estimation. While PSMODE achieves higher accuracy on normal data, its accuracy decline on soiled data is more pronounced when using fewer cameras. In contrast, SSMODE demonstrates greater robustness against soiled data with a reduced number of views. The qualitative results in Figure 13 illustrate that PSMODE predicts more detailed and accurate depth information, while SSMODE exhibits better performance on soiled data with only 2 views.
Moreover, compared with the results of existing methods in Table III, PSMODE and SSMODE achieve comparable performance with only 2 views. The experiments demonstrate that the proposed two frameworks are compatible with different numbers of views.
VI-F Ablation Study
We leverage spherical convolution in the feature extraction module and remove the image cropping during training PSMODE. We also add RGB panoramas and confidence maps into the depth fusion network. To verify the improvement of each component, we adopt ablation experiments on the two stages of SSMODE. Table VI shows the ablation studies of the omnidirectional stereo matching network. The results show that using panoramas without cropping and applying spherical convolution improve the performance. Table VII illustrates the ablation studies of the depth map fusion network. The results show that the fusion stage improves the quality of depth maps. The rows of the table gradually show the improvement of adding each component into the network.
TABLE V: Quantitative results of PSMODE and SSMODE with different view numbers on Deep360 dataset. The metrics refer to depth errors
@@ VI-F Ablation Study
TABLE VI: Ablation studies for omnidirectional stereo matching on Deep360. We compare the results of the proposed network with and without Input Image Cropping (Cr) and Spherical Convolution (SC). The metrics refer to disparity errors
@@
TABLE VII: Ablation studies for multi-view depth fusion in PSMODE on soiled Deep360. We compare the performance of the proposed fusion network with and without RGB images and Confidence maps. We list the result that without fusion (w.r.t the results of stereo matching stage in PSMODE) in the first row as the baseline. The metrics refer to depth errors
@@
VI-G Comparison of Two-stage and One-stage Methods
As illustrated in Table III and Table IV, PSMODE outperforms SSMODE on the Deep360 dataset when utilizing four cameras. According to the results in Table V the accuracy of PSMODE experiences a more significant decrease on soiled data when the number of cameras decreases. PSMODE fuses the results of pairwise stereo matching, which can integrate the information of different views to mitigate the distortion and blind points of the GEER projection. Consequently, the number of views has a more pronounced impact on PSMODE. On the other hand, SSMODE constructs a unified cost volume for all cameras and exhibits slightly lower accuracy compared to PSMODE. However, SSMODE demonstrates greater robustness to variations in the number of input cameras.
We also compare the video memory usage and time consumption of PSMODE and SSMODE, with the details provided in Table VIII. The two-stage PSMODE consists of an omnidirectional stereo matching network and a depth fusion network, and both networks can be trained independently. Therefore, PSMODE can employ a larger model with more video memory. However, the two-stage pipeline of PSMODE costs more time during the inference phase. SSMODE requires more video memory in training but has a faster inference speed. Moreover, PSMODE needs to estimate the depth map for each camera pair by stereo matching, which increases the computational complexity. As the number of cameras increases, the computational complexity of PSMODE grows significantly, resulting in reduced efficiency of the method.
@@ VI-G Comparison of Two-stage and One-stage Methods
TABLE VIII: Comparison of PSMODE and SSMODE in training memory and inference time. We use NVIDIA RTX3090 for training and inference, and set the resolution of input panoramas as 1024$\times$512 and batch size as one
@@
VII Discussion and Conclusion
VII-A GEER projection
As shown in Figure 2(a), we transform the panoramas into GEER projection to build the epipolar constraint for binocular 360 cameras and represent the disparity with the angle difference. However, for those points on the x-axis (line $O_{l}O_{r}$), the angle $\phi$ is always the same on left and right cameras:
@@ VII-B Conclusion
In practical applications, fisheye cameras are often more prevalent than 360 cameras[68]. We consider fisheye images as partially occluded spherical images. Thus, the proposed Generalized Epipolar Equirectangular (GEER) projection and depth estimation algorithms are applicable to this setting. However, fisheye cameras have smaller field-of-view (FoV) and exhibit limited overlapping areas between cameras when compared to 360 cameras. Notably, PSMODE requires a larger overlapping area since it relies on stereo matching to obtain initial depth maps. SSMODE also requires a common field of view for the cameras, and areas where only one camera is visible will lead to degraded monocular depth estimation. Consequently, the processing of overlapping and non-overlapping areas emerges as an open problem in multi-view omnidirectional depth estimation. Furthermore, we will study the real-time optimization of the MODE algorithms in future work to improve the efficiency and practical applicability.
Acknowledgments
This work was funded by the Suzhou Science and Technology Plan (Frontier Technology Research Project) SYG202334.

### v2

@@ 4.2.1 Pairwise Stereo Matching
4.2.2 Omnidirectional Depth Fusion
@@
We align the depth maps and confidence maps to the same viewpoint based on the extrinsic matrix and visibility. As shown in Fig. 5, the depth fusion network generally follows Unet[66], containing two encoders and one decoder. One encoder takes concatenation of the aligned depth maps and confidence maps as input to effectively aggregate the depth feature. and the other takes RGB panoramas as input to extract context and boundary features. Subsequently, these two types of features are fused through a multi-scale feature fusion block to generate the more comprehensive and informative feature maps. Finally, the decoder utilizes fused feature maps to perform regression and predict the final depth map.
@@
4.3 Spherical Sweeping Multi-view Omnidirectional Depth Estimation(SSMODE)
@@
As shown in Fig. 6, the proposed SSMODE first extracts features of input panoramas with the Spherical Feature Extraction Module. By employing the GEER projection, we establish the angular coordinate system $(\theta,\phi)$ and define a collection of hypothetical spheres at different depths $\rho_{i}$. The features for each point $(\rho_{i},\theta,\phi)$ across different views are obtained through the camera extrinsics, and these features are concatenated to construct the spherical cost volume. As illustrated in Fig. 3 and Equation(5), the unified spherical cost volume contains the matching cost of each pixel at each hypothetical depth. Similar to stereo matching, the spherical cost volume is also represented as a 5D tensor with a shape of $(B\times C\times D\times H\times W)$, where $(H\times W)$ denotes the angular coordinate of the sphere and $D$ represetns the number of hypothetical spheres. Subsequently, the 3D stacked hourglass module is employed to aggregate the multi-view spherical matching cost. Based on the regressed comprehensive matching cost, the weights of different depth are calculated, and the final depth is obtained by weighted summation:
@@
5 Dataset
Table 1: Overview of the proposed datasets and other published datasets
@@ 5 Dataset
As summarized in 2.3, although many datasets have been proposed for omnidirectional depth estimation, no 360 stereo dataset for outdoor road scenes is available due to the difficulty of acquiring 360 outdoor 3D datasets in the real world. Therefore, we create a public available 360 multi-view dataset Deep360 based on the CARLA autonomous driving simulator. Fig. 7 shows some examples of the dataset. We set four 360°cameras and arrange the cameras on a horizontal plane to form a square with side length as one meter, as shown in Fig. 1(a). The cameras are numbered from 1 to 4. Any two of the cameras can form a stereo pair, so there are 6 $(C_{4}^{2})$ pairs in total. Each frame consists of six pairs of rectified panoramas, which cover all the pairwise combinations of four 360 cameras, six corresponding disparity maps and one ground truth depth map. All these images and maps have a resolution of $1024\times 512$. To acquire realistic 360 outdoor road scenes with high variety, we make the car with 360 cameras in CARLA drive automatically [58] in six different towns and spawn many other random pedestrian and vehicles.
We also provide a soiled version of the Deep360 dataset, which can be used to train and evaluate 360 depth estimation algorithms under harsh circumstances in autonomous driving. Deep360-Soiled contains panoramas soiled or affected by three common outdoor factors: mud spots, water drops and glare, as illustrated in Fig. 1(e). An overview of the proposed dataset and other published 360 datasets is listed in Table 1.
6 Experiment Results
Table 2: Quantitative results of stereo matching on the proposed Deep360 dataset. The metrics refer to disparity errors
@@ 6 Experiment Results
6.1 Experiment Settings
6.1.1 Datasets
@@ 6.1.3 Metrics
6.2 Omnidirectional Stereo Matching
The existing binocular stereo matching algorithm is able to directly predict the spherical binocular disparity map at arbitrary relative positions based on the GEER projection method. Thus, we first evaluate the proposed omnidirectional stereo matching network on the Deep360 dataset and compare it with the excellent stereo matching algorithms PSMNet[20], AANet [28] and CREStereo[31], as well as the omnidirectional method 360SD-Net[5]. For these approaches, we use the pre-trained models from the authors and follow their hyperparameters to finetune on Deep360. Fig. 8 shows the qualitative results of omnidirectional stereo matching on the deep360 dataset. The quantitative results in Table 2 illustrate that our stereo matching network with spherical feature learning achieves SOTA performance on 360 stereo matching.
We also present the results of stereo matching of two 360 cameras at different relative positions in Fig. 9. As shown in Fig. 8 and Fig. 9, the proposed GEER projection establishes the epipolar constraint of binocular 360 cameras at arbitrary relative positions. The results show that the proposed stereo matching method with spherical feature extraction module achieves high precision with clear details.
6.3 Omnidirectional Depth Estimation
@@
Table 3: Quantitative results of omnidirectional depth estimation on the proposed Deep360 dataset. The metrics refer to depth errors
@@
As shown in Table 3, PSMODE and SSMODE perform favorably against SOTA omnidirectional depth estimation methods, especially on the dataset with soiled panoramas. We also compare the result of PSMODE with and without the fusion stage in Table 3. As the results show, the multi-view depth fusion stage significantly improves the accuracy of omnidirectional depth estimation. As demonstrated in Table 3 and Fig. 11, the accuracy degradation of the proposed methods on the soiled data is significantly lower than that of existing methods. The comparison demonstrates the robustness of the proposed multi-view depth estimation methods against camera soiling. We also evaluate the proposed methods on 3D60 dataset and illustrate the results in Table 4 and Fig. 10. The proposed PSMODE and SSMODE achieve high accuracy on both indoor and outdoor scenes. In this paper, we leverage all three stereo pairs within the 3D60 (left-right,up-down,up-right) in the depth fusion stage of PSMODE by employing the GEER projection. Thus, the results in Table 4 is better than those reported in the conference version[11].
Table 4: Quantitative results of omnidirectional depth estimation on 3D60 dataset. The metrics refer to depth errors
@@
6.4 Results on Real Scenes
We use the best PSMODE model trained on Deep360 to predict 360 depth maps on real-scene data. We use four Insta One X2 360 cameras to build the camera system, as shown in Fig. 1(b). Fig. 12 illustrates that the proposed algorithm also achieves an accurate depth estimation on real-scene data.
@@ 6.4 Results on Real Scenes
6.5 Evaluation of Different Numbers of Views
@@
To evaluate the performance of PSMODE and SSMODE under varying view conditions, we conducted experiments on the clear and soiled Deep360 dataset using different numbers of views (4, 3, 2). Table 5 and Fig. 13 indicate that as the number of views decreases, both PSMODE and SSMODE experience increase in error of depth estimation. While PSMODE achieves higher accuracy on normal data, its accuracy decline on soiled data is more pronounced when using fewer cameras. In contrast, SSMODE demonstrates greater robustness against soiled data with a reduced number of views. The qualitative results in Fig. 13 illustrate that PSMODE predicts more detailed and accurate depth information, while SSMODE exhibits better performance on soiled data with only 2 views.
Moreover, compared with the results of existing methods in Table 3, PSMODE and SSMODE achieve comparable performance with only 2 views. The experiments demonstrate that the proposed two frameworks are compatible with different numbers of views.
6.6 Ablation Study
We leverage spherical convolution in the feature extraction module and remove the image cropping during training PSMODE. We also add RGB panoramas and confidence maps into the depth fusion network. To verify the improvement of each component, we adopt ablation experiments on the two stages of SSMODE. Table 6 shows the ablation studies of the omnidirectional stereo matching network. The results show that using panoramas without cropping and applying spherical convolution improve the performance. Table 7 illustrates the ablation studies of the depth map fusion network. The results show that the fusion stage improves the quality of depth maps. The rows of the table gradually show the improvement of adding each component into the network.
Table 5: Quantitative results of PSMODE and SSMODE with different view numbers on Deep360 dataset. The metrics refer to depth errors
@@ 6.6 Ablation Study
Table 6: Ablation studies for omnidirectional stereo matching on Deep360. We compare the results of the proposed network with and without Input Image Cropping (Cr) and Spherical Convolution (SC). The metrics refer to disparity errors
@@
Table 7: Ablation studies for multi-view depth fusion in PSMODE on soiled Deep360. We compare the performance of the proposed fusion network with and without RGB images and Confidence maps. We list the result that without fusion (w.r.t the results of stereo matching stage in PSMODE) in the first row as the baseline. The metrics refer to depth errors
@@
6.7 Comparison of Two-stage and One-stage Methods
As illustrated in Table 3 and Table 4, PSMODE outperforms SSMODE on the Deep360 dataset when utilizing four cameras. According to the results in Table 5 the accuracy of PSMODE experiences a more significant decrease on soiled data when the number of cameras decreases. PSMODE fuses the results of pairwise stereo matching, which can integrate the information of different views to mitigate the distortion and blind points of the GEER projection. Consequently, the number of views has a more pronounced impact on PSMODE. On the other hand, SSMODE constructs a unified cost volume for all cameras and exhibits slightly lower accuracy compared to PSMODE. However, SSMODE demonstrates greater robustness to variations in the number of input cameras.
We also compare the video memory usage and time consumption of PSMODE and SSMODE, with the details provided in Table 8. The two-stage PSMODE consists of an omnidirectional stereo matching network and a depth fusion network, and both networks can be trained independently. Therefore, PSMODE can employ a larger model with more video memory. However, the two-stage pipeline of PSMODE costs more time during the inference phase. SSMODE requires more video memory in training but has a faster inference speed. Moreover, PSMODE needs to estimate the depth map for each camera pair by stereo matching, which increases the computational complexity. As the number of cameras increases, the computational complexity of PSMODE grows significantly, resulting in reduced efficiency of the method.
@@ 6.7 Comparison of Two-stage and One-stage Methods
Table 8: Comparison of PSMODE and SSMODE in training memory and inference time. We use NVIDIA RTX3090 for training and inference, and set the resolution of input panoramas as 1024$\times$512 and batch size as one
@@
7 Discussion and Conclusion
7.1 GEER projection
As shown in Fig. 2(a), we transform the panoramas into GEER projection to build the epipolar constraint for binocular 360 cameras and represent the disparity with the angle difference. However, for those points on the x-axis (line $O_{l}O_{r}$), the angle $\phi$ is always the same on left and right cameras:
@@ 7.2 Conclusion
In practical applications, fisheye cameras are often more prevalent than 360 cameras[70]. We consider fisheye images as partially occluded spherical images. Thus, the proposed Generalized Epipolar Equirectangular (GEER) projection and depth estimation algorithms are applicable to this setting. However, fisheye cameras have smaller field-of-view (FoV) and exhibit limited overlapping areas between cameras when compared to 360 cameras. Notably, PSMODE requires a larger overlapping area since it relies on stereo matching to obtain initial depth maps. SSMODE also requires a common field of view for the cameras, and areas where only one camera is visible will lead to degraded monocular depth estimation. Consequently, the processing of overlapping and non-overlapping areas emerges as an open problem in multi-view omnidirectional depth estimation. Furthermore, we will study the real-time optimization of the algorithms in future work to improve the efficiency of 3D measurement in practical.